OpenStudy (anonymous):

Choose the correct simplification of 9x2(4x + 2x2 – 1) EXPLAIN HOW TO DO IT??!?!??!??!?!??!

OpenStudy (lgbasallote):

well in my book, simplification is factoring...however, some say simplifying means expanding...so how were you taught?

OpenStudy (anonymous):

well here let me show you these are the answers A: 18x4 + 36x3 − 9x2 B: 18x4 − 36x3 + 9x2 C : 36x4 + 18x3 − 9x2 D: 36x4 − 13x3 + 9x2 So what do you think would get us here?

OpenStudy (lgbasallote):

ahh so you expand then

OpenStudy (lgbasallote):

just distribute 9x^2 into 4x + 2x^2 - 1 you know how right?

OpenStudy (anonymous):

yes multiply right?

OpenStudy (anonymous):

9x^2 * 4x = 36x^3 9x^2 * 2x^2 = 18x^4 9x^2 * -1 = - 9x^2

OpenStudy (anonymous):

right?

OpenStudy (lgbasallote):

what you wrote are right...now add those

OpenStudy (anonymous):

add them ?? why??

OpenStudy (lgbasallote):

well....because that's distributive property.. 2x (3x + 1) 2x * 3x = 6x^2 2x * 1 = 2x add them 6x^2 + 2x forgot?

OpenStudy (anonymous):

so is that answer?

OpenStudy (lgbasallote):

1) look at the problem i wrote 2) does that answer resemble any of your choices?

OpenStudy (anonymous):

nooo???

OpenStudy (lgbasallote):

like i said.. what you wrote is right already. just add them together to get the final answer

OpenStudy (anonymous):

but it says its wrong on the test ? :?

OpenStudy (radar):

normally you would arrange the powers of x in a descending order. From your results you would start with:\[18x ^{4}\]

OpenStudy (radar):

Then follow with:\[18x ^{4}+36x ^{3}\] and so on.

OpenStudy (radar):

Did you say choice Choice A. was wrong?